OpenStudy (anonymous):

A satellite in a nearly circular orbit is 2000 km above earth's surface.The raidus of earth is approx. 6400 km. If the satellite completes its orbit in 12 hours, calculate the speed of the satellite in kilometers per hour. then convert that speed to meters per second

OpenStudy (anonymous):

Do you know any formulas for circular motion? Identify the given informatioon and see if you have a formula in which only speed is missing.

OpenStudy (anonymous):

orbital velocity v= 2*3.14*a/T = 6.28*8400/12 = 4396 km/hr =2.75x10^3 m/s